OpenStudy (anonymous):

can somebody please answer this for me and show me how? i really dont understand! ^12sqrtx^5*^6sqrtx

OpenStudy (anonymous):

somebody imm begging pleasreeee!:(

OpenStudy (anonymous):

i dont understand '^12' this is 'to power 12' - what to the power 12?

OpenStudy (vijay):

type your question again, with proper symbols

OpenStudy (anonymous):

\[\sqrt[12]{x}\times \sqrt[6]{x}\] this is what it looks like

OpenStudy (dumbcow):

\[\sqrt[12]{x^{5}}*\sqrt[6]{x} = x^{5/12}*x^{1/6} = x^{5/12+1/6} = x^{7/12}\]

OpenStudy (anonymous):

yup